Suppose because of anticipated higher inflation in the near future, the Fed wants to decrease country’s money supply by $100 billion now.

a)   If the reserve ratio R = 10% and the Fed wants to use its open market operations policy tool, will it buy or sell U. S. government bonds?

b)   What would be the amount of bonds the Fed will buy or sell in the market? Show your calculations.

1
Expert's answer
2021-05-05T13:33:58-0400

Given,

Money supply = $100 billion


(a)

Open market operation refers to the selling and buying of government bonds by the central bank. Open market sale is a way in which the fed sells government bond or securities due to which money supply contracts in the economy by reducing the money holding in hand of public. So, in order to reduce money supply, the Fed must sell U.S government bonds.


(b)

R = 10%

Money multiplier can be computed as:

"money" "multiplier=\frac{1}{R}"


"money" "multiplier=\frac{1}{10}"


In order to decrease money supply by $100 billion, the Fed should sell U.S government bonds worth:

"\$100" "billion" "= money" "multiplier" "\times Worth" "of" "U.S" "bond"

"\$100" "billion""=10 \times""Worth" "of" "U.S" "bond"


"Worth" "of" "U.S" "= \$10" "billion"